# Collective Intelligence Outperforms Individual Talent: A Case Study in League of Legends

### v1

@@ 4.2 Collective Intelligence Outperforms Individual Intelligence
Furthermore, k-means clustering is used to profile teams and individual players based on these factor scores. The elbow method identified the optimal number of clusters for both individuals and teams to be $N=3$. Such clusters for individual players and teams is further labeled based on the average distribution of performance factors in each cluster. That is, individual player clusters are labeled as acquiring, sharing and average, while team clusters are labeled as cooperative, non-cooperative and average. From the experimental dataset containing 31,055 matches, there is a total of 310,550 players (10 players in each match) and a total of 62,110 teams (2 teams in each match). Table 3 shows the distribution of winning rate across different clusters of teams and individual players. Additionally, we perform cross-tabulation analysis on the distribution of winning rate in different team clusters having a majority of different types of players in Table 4. Table 5 presents a consolidated view of Table  3 and Table 4 comparing the winning rate of cooperative teams having a majority of average or sharing players (i.e. medium to low-performing individuals) with non-cooperative teams having a majority of average or acquiring players (i.e. medium to high-performing individuals).
@@ 4.3 Highly-skilled players exhibit higher levels of collective intelligence
[…]
Figure 5: Higher ranking teams (Diamond and above) exhibit higher levels of collective intelligence signified by lower values of effective graph resistance, in-degree centrality and out-degree centrality, compared to lower ranking teams (Silver and below).
@@ 5 Discussion
In this paper, we argue and present evidence for the statement – collective intelligence outperforms individual talent via a case study in League of Legends. By analyzing the frequency of wins, the results show that acting on the collective interest of a group leads to a higher likelihood of winning games rather than relying on individual skills and abilities. Moreover, we demonstrate that cooperative teams with low performing individuals outperform non-cooperative teams with high performing individuals.
Previous work by Mora-Cantallops and Sicilia report that centralization of resources is detrimental to team efficiency and performance [24], and our findings align with it. However, unlike this previous work in which the graph metrics – team in-degree centrality and team out-degree centrality were computed based on kills assistance, our work accounts for all forms of assists (in securing game objectives) including a tailored measure of passive assistance in positional map-pressure. Likewise, this work is the first study in which effective graph resistance is modeled to study team networks in League of Legends. Another work by Kim et. al [19] analyze group collective intelligence in controlled participants, however this work incorporates a large-scale participation of anonymous, more representative set of players, thereby enhancing the validity and general applicability of the findings. In their work, they show that higher tiered teams exhibit higher levels of collective intelligence which is congruent with our findings where we present statistically significant differences in the levels of collective intelligence across teams of different tier rankings. Additional concerns regarding the validity of our work is clarified in Appendix.
@@
The phenomenon of collective intelligence in multiplayer online gaming environments merits investigation into diverse gaming genres, for instance in First-Person Shooter (FPS) games, Massive Multi-player Online Role Playing Games (MMORPGs). Unlike FPS games (which mostly focus on individual reflex and mechanics) and MMORPGs (in which games can span over a longer time period), the authors chose to study collective intelligence in MOBA games due to their well-defined team structures, player roles, resource and objective trade-offs that bear similarities to collective actions in real-world problems. Specifically, the authors chose to conduct this study in League of Legends because – (a) it is the most popular MOBA game played worldwide, (b) features API endpoints that allow access to large-scale structured data, and (c) personally due to the authors’ familiarity and inclination towards the game. The framework that is presented in this work can also be used in other MOBA games like DOTA 2 to quantify the levels of outperformance. However, implementation will require tweaking the source code (data structures and API calls).
On a broader perspective, our results present the contradictory dynamics observed in the behavior of an individual when optimizing for self vs when optimizing for the collective. Such dynamic behaviors can also be observed in the design of sustainable infrastructure products. For instance, individuals can be tempted to use non-sustainable products which are inexpensive (immediate lower negative payoff) compared to eco-friendly products which can be costly (higher negative payoff) to implement but over a long run offers a greater benefit to the entire population [29]. Insights presented in this study can thus stipulate independent individuals to prioritize collective altruistic behaviors over individual predisposition in complex social systems. Furthermore, collective efforts can be crucial to attain the desired goals not just within online gaming ecosystems, but also in other more fundamentally critical aspects of our online (as well as offline) existence. For instance, collective efforts can help restrict the misinformation that is rampant in online social networks. Policies and regulations set up by the decentralized and collective efforts of the population in socio-economic systems such as consumer markets, governments, etc. prevent monopoly and abuse of authority resulting in products and systems that are favorable for the whole collective.
@@ 6 Conclusion
In this paper, we explored the presence of collective intelligence among players in the MOBA game, League of Legends. Our findings conclude that collective intelligence characterized by sharing of resources, flow of information and cooperative behavior outperforms individual intelligence. Our results reiterate the paradoxical imperative for collaboration and cooperation in highly competitive environments for achieving optimal outcomes. Further work complementing this research can be done by testing for the performance of collective intelligence and individual intelligence in real-world applications beyond gaming environments.
@@ Acknowledgements
Data and Code Availability
@@
Relevant source code and sample dataset used in the analyses outlined in the paper is available at https://github.com/TDI-Lab/Collective-Intelligence-LoL.
Appendix A Appendix
A.1 League of Legends Gameplay Basics
There are different modes of gameplay available in League of Legends, for instance – ARAM, URF, Summoner’s Rift, Team Fight Tactics, etc. Ranked matches in Summoner’s Rift are the most popular and competitive modes of gameplay within the game in which two teams (each comprising of 5 players) compete to win against the other by destroying the opponent team’s base structure called nexus. Throughout the game, different types of resources and objectives are spawned, which the two teams compete for via fighting in order to earn gold and craft items which ultimately helps in destroying the nexus. Figure 6 shows a prototype map of summoner’s rift gameplay. The map comprises of three lanes – top-lane, mid-lane and bottom-lane. Accordingly, a player is assigned the role for each of the lanes. The remaining two players are assigned the roles of jungle and support. The jungler has the responsibility of moving around the lanes and assisting teammates to gain further gold via opponent kills and/or securing in-game objectives and resources. The support player accompanies the bottom-lane player (often called the Attack-Damage-Carry) providing support and utility for the player, so that over the duration of the game, the bottom-laner can scale and later ‘carry’ the game.
Figure 6: The three lanes, 2 jungles and their subdivisions. Small red and blue circles denote tower structures for the respective teams. The darker, inner-circle segment in the top-right and bottom-left corners denote each teams’ Nexus
A.2 Champion Basics
In games, each player steers independent control on in-game avatar called a champion. League of Legends features 169 unique playable champions, each with their own unique set of abilities. Such abilities can be used to deal damage, heal, shield or cast other effects to enemy or ally champions. Each champion is also equipped with 2 set of summoner spells, which provide various effects and utilities such as healing, shielding, slowing, transporting, etc. A champion requires Health Points to stay alive. A dead champion is re-spawned after a certain interval of time. Champions require mana points to cast their abilities. Items are purchased using gold. As champions gain experience and level-up, the magnitude of their abilities and base statistics rise. Figure 7 shows an in-game champion with its set of abilities, spells and item crafting.
Figure 7: Details of an in-game champion
A.3 Resources: Gold, Experience and Vision
Experience is a metric that enables players to increase their Champion’s level, in turn, increasing the Champion’s strength. Experience is primarily gained by farming minions or jungle camps. However, it is not imperative that a player kills the minion to be awarded experience. Gold is used by players to purchase in-game items, making their champion stronger. Gold is one of the most important factors which contributes towards winning the game. Gaining gold and buying items enables a champion to carry their lead forward by helping their team achieve additional objectives. Gold is earned in various ways – passively at a steady rate of time, scoring kills on enemy champions, last-hit on minions and elite-monster kills, destroying enemy tower structures, etc. Vision is a valuable resource measured by ‘vision score’. However, unlike gold or experience, no advantage is given to the player in terms of strength or currency. Players spend their gold to purchase wards – these items are placed strategically on the map to reveal dark areas that cannot be seen otherwise. These dark areas may reveal the position of an enemy player, thus, yielding a strategic advantage to the other team. Moreover, players can destroy wards placed by the opposing team – denying them such information. Doing these actions (purchasing, placing and destroying wards) increases their vision score.
A.4 Turrets and Elite-Monsters
Turrets (also known as towers), are positioned in every lane and surround each teams’ base. Turrets will attack all enemy characters within their range, one at a time. Players gain gold by attacking and destroying the opposing teams’ turrets. Destruction of an enemy turret is an example of achieving an objective. Turrets can also be destroyed by minions, in which case players receive no monetary benefit. Elite monsters are considered common objectives as they do not belong to a team, spawning mid-way at the rift. All players in the game look to kill them and reap their benefits. A small amount of gold is rewarded for securing elite monsters objectives. However, their main benefit is improving the base statistics of players (of the team that secure the objective) such as movement speed, healing rate, abilities cool-down, etc, which may be temporary or over the duration of the entire game.
A.5 Kills and Assists
A player is awarded with a kill score if the player uses their abilities or spells such that an enemy’s health points is reduced to zero. A player scoring a kill on enemy champion is rewarded with a significant amount of gold. A player scores an assist when the player helps a team-mate score a kill on an enemy champion. Assists are also awarded for tower destruction and elite-monster kills. Assistance can be in the form of dealing damage to enemy champions or providing utility effects such as healing, shielding, etc. to ally champions that results in the kill. Players receive a small amount of gold when providing assists.
A.6 Positional map-pressure assist
A new tailored metric called the “positional map-pressure" assist, which reflects how a player within a team exhibits intelligent map positioning thereby resulting in an opponent kill, is computed for each match. Figure 8 depicts how positional map-pressure can occur in-game. Opponent kill assists, tower destruction assists or elite monsters kill assists by a given player can be directly retrieved by processing the match timeline data. However, positional map-pressure assist is not directly accounted within the game. As such, by using the match timeline data, positioning of all team players an opponent kill event were identified, and a player was awarded a positional map-pressure assist if the player’s position was within a specified radius during the kill event.
Figure 8: An instance of gameplay on the Summoner’s Rift map in which an opponent kill is scored due to positional map-pressure. Blue team player A is attempting to score an opponent kill on red team player Y, while Y attempts to flee. Another blue team player B obstructs Y from escaping, thus inciting Y to engage in a fight resulting in a kill score for blue. While B does not actively assist A in the kill process, B’s positioning induces map pressure causing Y’s death. B is awarded map-pressure assist if B is positioned within a specified radius during the kill event.
A.7 Ranking System and League Points
The ranking system used in League of Legends employs a two-tiered structured ladder designed to measure and reward player skill through competitive gameplay in ranked matches. The ranking system consists of ten tiers: Iron, Bronze, Silver, Gold, Platinum, Emerald, Diamond, Master, Grandmaster, and Challenger. Each tier below Master is subdivided into four divisions (for example, Gold IV to Gold I), with IV being the lowest and I the highest within a tier.
League Points (LP) are the core mechanism that enable progression in the ranking system. Players earn LP by winning matches and lose LP upon defeat. In each sub-division of tiers below Master, LP are ranged between 0 and 100, such that upon exceeding 100 LP for a given sub-division within a tier, the player is promoted to the next immediate division (or the next tier if current division is I) and LP is reset to 0. For promotion in tiers above Master, there is no fixed rule, but rather depends on pool size of players in the upper tiers as well as higher amount of LP is required. The amount of LP gained or lost per match is influenced by several factors, including a hidden matchmaking Rating (MMR), which reflects a player’s skill level relative to their rank. Higher MMR results in larger LP gains for victories and smaller losses for defeats
A.8 Data Retrieval Process
Data on League of Legends ranked matches played on any of the official server maintained by Riot Games is freely accessible via a set of RESTful API endpoints. Figure 9 shows the overall data-retrieval process used in the study. An initial set of 10 (one for each ranking tier) random, hand-picked summoner names are stored in a text file as seed data. A python script reads this seed configuration file and makes the following API calls –
  • summonerName endpoint: This API endpoint fetches the corresponding unique identifier, called the puuid, for a given summoner name.
  • puuid/ids endpoint: This API endpoint returns a list of match identifiers for the most recent N-games (configurable) played by a player with a given puuid.
  • matchid endpoint: This API endpoint fetches the details of a given match id for each player in both teams such as - gold earned, items crafted, experience gained, kills scored, champion roles and levels, etc. Details of each match is saved as a JSON flatfile.
  • matchid/timeline endpoint: This API endpoint fetches the details of various game event (such as kills, level-ups, items purchase, etc.) as it occurs chronologically over the duration of the game with a given match id. Timeline details of each match id is saved as a JSON flatfile.
  • entries/by-summoner/summonerId endpoint: This API endpoint retrieves the rank information data for a given summoner identified by their unique puuid.
The client script thus reads the initial seed summoner names and writes match data and match timeline files in the local directory. Such files are merged and data filters, concerning game-type, duration, disconnection, etc. are applied to remove inconsistent rows.
Figure 9: Data retrieval process via RESTful APIs provided by Riot Games
A.9 Threats to Validity
There may be concerns of initial selection bias when handpicking the seed summonerNames used in the data collection process. However, we ensure that the additional fetch retrieves data for players that are not in the initial list. In 50% of the dataset, it is ensured that a single unique player has played no more than 2 games, resulting in a total of 164K unique players. Thus, any selection bias over the initially selected 10 seed summoners is unlikely to impact the results. Additionally, the collected dataset features games played exclusively within the European servers. League of Legends is played on different servers across different geographical locations i.e., America, Europe, Asia and Oceania, with notable differences in players’ pool size and skills. However, we expect data from different servers to exhibit similar pattern of results. In-game metrics such as towers destroyed, elite monster kills, enemy champion kills and minion kills are highly correlated with gold earned. Thus, these features are removed. Variance inflation test yields values of $VIF>5$ for the feature gold earned, which signifies high correlation with other features. However, it was not removed from the feature set as it is the most contributing feature for the acquiring factor. The suitability of factor analysis at the individual and collective levels yield KMO values of $0.57$ and $0.61$ respectively, which is poor but acceptable. Our results demonstrate that the winning rate for a cooperative team with a majority of resource acquiring players is as high as 94%. However this may be misinterpreted. It does not mean that the same cooperative team with the same set of high-performing individuals will win 94% of their games. This is because the Riot Games’ matchmaking algorithm ensures that as players win games, they are positioned higher in the ranking ladder, as a result they are matched with relatively more-skilled, difficult opponents when winning, and ensuring that most players roughly win 50% of their games. Finally, since the outcome of a match is a zero-sum game (no draws are applicable for a match), a team can lose not because it is bad, but the opponent has a small advantage over the team. On the other hand, a non-cooperative team can also win exploiting a blunder from the opponents. Such extreme cases have not been handled in the dataset.

### v2

@@ 4.2 Collective Intelligence Outperforms Individual Intelligence
Furthermore, k-means clustering is used to profile teams and individual players based on these factor scores. The elbow method identified the optimal number of clusters for both individuals and teams to be $N=3$. Such clusters for individual players and teams is further labeled based on the average distribution of performance factors in each cluster. That is, individual player clusters are labeled as acquiring, sharing and average, while team clusters are labeled as cooperative, non-cooperative and average. From the experimental dataset containing 31,055 matches, there is a total of 310,550 players (10 players in each match) and a total of 62,110 teams (2 teams in each match). Table 4 shows the distribution of winning rate across different clusters of teams and individual players. Additionally, we perform cross-tabulation analysis on the distribution of winning rate in different team clusters having a majority of different types of players in Table 5. Table 6 presents a consolidated view of Table  4 and Table 5 comparing the winning rate of cooperative teams having a majority of average or sharing players (i.e. medium to low-performing individuals) with non-cooperative teams having a majority of average or acquiring players (i.e. medium to high-performing individuals).
@@
Last but not least, a visual approach is also used to demonstrate the superior performance of collective intelligence in Figure 6. For this, we pick up one of the strongest indicators of collective intelligence in Table 2, i.e. the effective graph resistance, to rank the teams and their interactions. The top-5 and bottom-5 teams are presented to visually inspect the nature of the interactions. As shown in Figure 6, the top-5 (lowest EGR) teams show diverse interactions between the players and most of those win the match. In contrast, the bottom-5 (highest EGR) teams have more limited and isolated interactions between the players and all of those result in a loss.
Figure 6: Chord diagrams based on data of graph networks are used here as a visual approach for the communication and analytics of collective intelligence within League of Legends. The team interactions between the players are visualized for the ones with the top-5 and bottom-5 values of effective graph resistance (ranked from lowest to highest). For each team, the result of match being a loss or win is indicated. There are five nodes in each chord plot signifying unique player positions in the game. The thickness of the chords signify the weight of successful interactions (assistance) between players in a team relative to the overall performance of the team. Teams with top-ranked EGR (lower) values exhibit multi-directional flow of interactions among players in a team while teams with bottom-ranked EGR (higher) values exhibit minimum interactions among each other.
@@ 4.3 Highly-skilled players exhibit higher levels of collective intelligence
To test for H2, we investigated the association of graph-based collective intelligence metrics with the average rank of a team. Players who excel in the game are signified by a higher positioning in the ranking ladder. We find that teams containing highly-skilled players, based on their average team rank, exhibit moderately higher levels of collective intelligence (signified by lower values of effective graph resistance) compared to less-skilled teams. When teams are partitioned into two groups i.e highly-skilled and less-skilled, based on the values of average team rank sampled at size N=8000 teams with even distribution of wins and loss for 5 random samples, statistically significant differences emerge between the groups with respect to collective intelligence indicators. Specifically, the distributions of effective graph resistance, in-degree centrality, and out-degree centrality differ significantly across the two groups (two-tailed t-test, p < 0.05 [32], see Figure 7). However, less-skilled teams exhibit lower degrees of centrality compared to highly-skilled teams. Such can be the case as highly-skilled teams are more aware of player positions and roles, and thus tend to focus resources towards specific players that can win the game.
Figure 7: Teams in higher ranks (Diamond and above) demonstrate greater collective intelligence for winning teams, as indicated by lower effective graph resistance, but higher in-degree and out-degree centrality, compared to lower-ranked teams (Silver and below). Randomly sampled winner and loser teams (5 times), sample size = 8000
@@ 5 Discussion
This paper unravels collective intelligence in League of Legends. The first hypothesis, H1:collective intelligence outperforms individual intelligence has been analyzed and demonstrated using classification models, clustering algorithms and cross-tabulation analysis. Thus, teams prioritizing collective goals win more often than those relying on individual skill. Even low-performing players in cooperative teams outperform high-performing individuals in noncooperative ones. Although League of Legends is designed as a collaborative team-based game, individuals are often matched up with random players which often shifts the emphasis from required coordinated team play to individual performance. The novel contribution of this work lies in the visual representation and analysis of team networks to re-iterate the importance of collective efforts over isolated individual skills. Visual analysis highlights complex interactions and low effective graph resistance as key indicators of collective intelligence. We are the first to use effective graph resistance for modeling team networks in League of Legends, offering new insights through visual analytics and communication. Building on Mora-Cantallops and Sicilia [28], who found resource centralization harms performance, our study extends this by incorporating all assist types, including positional map-pressure. Compared to the controlled settings of Kim et al. [23], our large-scale dataset of anonymous players provides broader validity.
For the second hypothesis, H2: highly-skilled players exhibit higher levels of collective intelligence compared to less-skilled players, we presented statistically significant evidence showing that highly-skilled teams and players (signified by the ranking system) exhibit higher levels of collective intelligence compared to less-skilled players via two-tailed t-test.
@@
The phenomenon of collective intelligence in multiplayer online gaming environments merits investigation into diverse gaming genres, for instance in First-Person Shooter (FPS) games, Massive Multi-player Online Role Playing Games (MMORPGs). Unlike FPS games (which mostly focus on individual reflex and mechanics) and MMORPGs (in which games can span over a longer time period), the authors chose to study collective intelligence in MOBA games due to their well-defined team structures, player roles, resource and objective trade-offs that bear similarities to collective actions in real-world problems. Specifically, the authors chose to conduct this study in League of Legends because – (a) it is the most popular MOBA game played worldwide, and (b) features API endpoints that allow access to large-scale structured data. The framework that is presented in this work can also be used in other MOBA games like DOTA 2 to quantify the levels of outperformance. However, implementation will require tweaking the source code.
Our results highlight the contrast between individual behavior when optimizing for self versus for the collective. This dynamic also appears in sustainable infrastructure, where individuals may choose cheaper non-sustainable products over costly eco-friendly ones that benefit society long-term [33]. The study suggests individuals may be guided to prioritize collective altruism in complex social systems. Such collective behavior is vital not only in online games but also in real-world contexts. For instance, it can help combat misinformation in online networks and support fairer outcomes in socio-economic systems like consumer markets and governments by preventing monopoly and power abuse.
@@ 6 Conclusion
In this paper, we explored the presence of collective intelligence among players in the MOBA game, League of Legends using quantitative methods and a visual approach to knowledge discovery. Our findings conclude that collective intelligence characterized by sharing of resources, flow of information and cooperative behavior outperforms individual intelligence. Our results reiterate the paradoxical imperative for collaboration and cooperation in highly competitive environments for achieving optimal outcomes. Further work complementing this research can be done by testing for the performance of collective intelligence and individual intelligence in real-world applications beyond gaming environments.
@@ Acknowledgements
Code, Dataset and Supplementary Materials
@@
Relevant source code used in the analyses outlined in the paper is available at https://github.com/TDI-Lab/Collective-Intelligence-LoL. Supplementary materials on game details and collected dataset is available under the corresponding Zenodo record [32].